Term # | Blocks added | Total Blocks
1 | 25 | 25
2. | 18. | 43
3. | 18. | 51
4. | 18. | 69
5. | 18. | 87
Nth. |. ? |. ?

Answer by Edwin McCravy(20056)

Term # | Blocks | Total 
       | added  | Blocks
------------------------------
  1.   |   25   | 25
  2.   |   18   | 43 = 25+1(18) 
  3.   |   18   | 51 = 25+2(18)
  4.   |   18   | 69 = 25+3(18) 
  5.   |   18   | 87 = 25+4(18)
  ...  | ...    | ...
  Nth. |   18   | 25+(n-1)18

Then simplify 25+(n-1)18
              25+18n-18
              7+18n

Edwin